# RAG-Driver: Generalisable Driving Explanations with Retrieval-Augmented In-Context Learning in Multi-Modal Large Language Model

### v1

@@ Abstract
Robots powered by ‘blackbox’ models need to provide human-understandable explanations which we can trust. Hence, explainability plays a critical role in trustworthy autonomous decision-making to foster transparency and acceptance among end users, especially in complex autonomous driving. Recent advancements in Multi-Modal Large Language models (MLLMs) have shown promising potential in enhancing the explainability as a driving agent by producing control predictions along with natural language explanations. However, severe data scarcity due to expensive annotation costs and significant domain gaps between different datasets makes the development of a robust and generalisable system an extremely challenging task. Moreover, the prohibitively expensive training requirements of MLLM and the unsolved problem of catastrophic forgetting further limit their generalisability post-deployment. To address these challenges, we present RAG-Driver, a novel retrieval-augmented multi-modal large language model that leverages in-context learning for high-performance, explainable, and generalisable autonomous driving. By grounding in retrieved expert demonstration, we empirically validate that RAG-Driver achieves state-of-the-art performance in producing driving action explanations, justifications, and control signal prediction. More importantly, it exhibits exceptional zero-shot generalisation capabilities to unseen environments without further training endeavours [^1] Project Page: https://yuanjianhao508.github.io/RAG-Driver.
@@ I Introduction
Driven by the emerging development of deep learning, autonomous driving has observed a paradigm shift from rules-based decision systems [66, 21] to data-driven learning-based approaches [28, 6, 36]. However, this comes at the cost of transparency in decision-making, especially for end-to-end autonomous driving systems which are considered black-box in nature [13]. Thus, in addition to precision in action control, explanation provision is key in ensuring trustworthy decision-making to reconcile the system’s decisions with end-user expectations to foster confidence and acceptance [79, 8, 57] in dynamic driving environments.
@@
Traditional approaches have mainly relied on attention visualisation [5, 7, 55] as a proxy to rationalise the decisions of the black-box systems or auxiliary intermediate tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] provide meaningful intermediate representation for decision-making. However, these methods do not engage end-users in the dialogue as they are one-directional and not readily comprehensible by the general users for the purpose of fostering trust and confidence. An alternative promising approach is the integration of natural language explanations [38, 33, 54], in particular through Multi-Modal Large Language Models (MLLMs) [1, 70]. These models, pre-trained on extensive web-scale datasets, demonstrate remarkable reasoning capacity, enabling the transformation of complex vehicular decision-making processes into more understandable narrative formats, thereby offering a new layer of explainability to conventional systems.
@@
While several early attempts have demonstrated the potential of MLLMs as general explainable driving agents [78, 76, 51], these methods fall short of human-level understanding. One of the limitations is their failure to generalise to unseen environments. A primary obstacle is the lack of high-quality annotated data [56], coupled with the significant domain shift across various datasets [23], which hinders the models’ generalisation capacity to novel environments outside of the training data distribution. Another critical challenge is the prohibitively expensive training requirement and the unsolved problem of catastrophic forgetting [39], which make re-training or fine-tuning impractical solutions due to the immense computational demands and severe performance degradation. Consequently, this further limits the models’ generalisability after deployment, as they struggle to effectively utilise new data in constantly evolving environments and driving scenarios.
@@
To address these challenges, we introduce RAG-Driver, a novel retrieval-augmented multi-modal large language model tailored for generalisable and explainable end-to-end driving. As illustrated in Fig.1, it outputs natural language texts corresponding to (1) the driving action and (2) justification of that driving action along with (3) numerical control signals based on driving videos. The natural language texts are aligned with control signals during in-context learning [10] to enable faithful introspective explanation provision. The novelty of RAG-Driver is the integration of retrieval-augmented in-context learning (RA-ICL) mechanisms that significantly improve generalisation performance in unseen driving environments. It allows efficient recall of similar driving scenarios as contextual information augmenting MLLM prediction through implicit meta-optimisation (Sec. III-C). Through extensive experiments, we show that RAG-Driver outperforms existing methods on both in-domain deployments as well as deployment in unseen environments (without any fine-tuning). By being grounded in analogical demonstrations, our framework significantly reduces the need for continuous retraining while enhancing the generalisability and quality of generated explanatory texts. Our primary contributions are as follows:
@@
    Achieving state-of-the-art introspective driving explanation performance on the standard benchmark BDD-X [38].
@@
    Demonstrating exceptional zero-shot generalisation capacity to unseen scenarios without training effort through a customised dataset Spoken-SAX featuring video sequences narrated by a professional driving instructor.
@@ II-A Explainable End-to-End Autonomous Driving
End-to-end learned driving [13] maps directly from raw sensor input to vehicle control signals. This data-driven, joint optimisation of perception, prediction, and planning can be simple and efficient [28]. In this area, various learning-based approaches, including behavior cloning [6, 12, 61, 83, 60], inverse optimal control [81, 65, 75], and reinforcement learning [36, 71, 11, 44] are promising. A key area of focus in this area is explainability [80], which is crucial for improving transparency and building trust towards wider public acceptance of autonomous systems [54, 22, 57]. One line of works leverages attention visualisation – either to directly identify salient regions of input images that are important for driving decision-making [37, 5, 7] or to assist feature aggregation for downstream motion planning tasks [55, 15, 77, 63]. Another line of work uses intermediate auxiliary tasks such as semantic segmentation [25, 32], object detection [16, 31], and affordance prediction [68, 45] which help to decoder latent representations to human-understandable representation. While these methods provide explainable mechanisms by associating decision-making processes with semantic or visual representations, they are not readily comprehensible by the general users for the purpose of fostering trust and confidence.
Alternatively, recent research shows promise in utilising natural language explanation. Several works develop specialist explainers [38, 41] using attention alignment in visual input and textual generation for grounded driving action explanation. ADAPT [33] uses a vision-language transformer with separate decoder for caption generation alongside control signal prediction. More recently, several works explore the potential of Multi-modal Large Language Models (MLLMs). Works such as, DriveGPT4 [78], Lingo [54], and DrivingMLM [76] have shown promising potential in general question-answering for driving and action planning. However, a common obstacle in both specialist and MLLM-based generalist models is the scarcity of data due to expensive annotation costs and significant domain gaps between different datasets, making the development of a robust and generalisable model an extremely challenging task. In our work, we aim to overcome these obstacles by employing a more robust inference paradigm of retrieval-augmented in-context learning to bridge domain gaps and circumvent the need for annotations in new domains.
@@ II-B Multi-Modal Large Language Model
Recent advancements in Large Language Models (LLMs) have paved the way for the emergence of Multi-modal Large Language Models (MLLMs) [1, 70]. Benefiting from scalable transformer-based architecture and web-scale training data, these models have demonstrated notable capabilities in general-purpose visual understanding tasks. One line of work focuses on modality fusion in the latent space, offering a scalable, end-to-end solution for MLLMs. For instance, Flamingo [3] and BLIP2 [43] fuse visual tokens to frozen LLM through gated attention and query transformers, respectively. LLaVA [48] and MiniGPT4 [86] use simple Multi-Layer Perceptron (MLP) with visual instruction tuning to align the pre-trained image encoder to LLM. Most relevant to us is the line of work focusing on video-language models such as Video-LLaVA [46] and Video-LLaMA [82], which integrate pre-trained video encoders into LLMs using strategies similar to those in image-based models.
With remarkable perception and reasoning capacity, MLLMs reveal promising potential in various robotics tasks such as reasoning [20, 2, 30] and planning [62, 9, 47, 35]. The most similar to us is the idea that engages a generalist foundation model for end-to-end embodied agent. PaLM-e [20] injects images, state estimates, and other sensor modalities to LLM and autoregressively produces natural language commands. RT-2 [9] and RT-X [58] fine-tune on image and low-level robot control signal pairs to perform end-to-end robot control. Specifically in driving, numerous approaches leverage language-only LLM for decision-making [79] and then augment it with external perception module feedback [53, 24, 14], designed chain-of-thought reasoning template [53, 52] or downstream planner [67, 69, 14] to form a system-level driving agent. Another more relevant line of work is end-to-end driving agents. DriveGPT4 [78] leverages a finetuned video-language model Valley [50] on driving specific visual instruction tuning based on BDD-X [38]. Dolphins [51] further use the designed Grounded chain-of-thought to enhance reasoning capacity. DrivingMLM [76] and Reason2Drive [56] scale up the driving visual instruction tuning dataset through simulator-based data engine and annotation of existing large-scale datasets, respectively. While these approaches have demonstrated the potentials of MLLM, the prohibitively expensive training cost and the unsolved problem of catastrophic forgetting, which makes re-training or fine-tuning post-deployment challenging, further limit their generalisation capacity to unseen driving environments. To solve this problem, we leverage a training-free retrieval-augmented in-context learning mechanism.
@@ II-C In-Context Learning and Retrieval-Augmented Generation
While LLMs demonstrate strong generative and reasoning capacity, there are still several issues associated with their output, such as hallucination [29], and slow knowledge updates [34]. In-context Learning (ICL) [10, 18] has emerged as a promising approach in LLM inference, potentially addressing several of these issues. This paradigm involves providing a test query alongside a few demonstration examples as contextual information. The LLM then generates an output for the test instance based on analogies drawn from context, without any updates to its parameters [64]. While ICL has been observed to enhance generalisability in various Natural Language Processing (NLP) tasks, its application in multi-modal contexts remains less explored, potentially due to the challenges associated with curating structured high-quality multi-modal ICL datasets. Retrieval-Augmented Generation (RAG) [42] is another important inference paradigm for LLM. It provides an external knowledge database to augment the compressed model knowledge within LLM in inference by dynamically retrieving relevant information pieces as contextual information. One of its promising applications is leveraging a systematic approach to curate In-Context Learning (ICL) examples. In this work, we build upon these inference paradigms and extend their application to Multimodal Large Language Models (MLLMs). We introduce a retrieval-augmented in-context learning mechanism through a curated multi-modal driving in-context instruction tuning dataset and a vector similarity-based retrieval engine specifically tailored for driving applications.
@@ III-A Multi-modal Large Language Model Architecture
Following the successful MLLM paradigm of Video-LLaVA [46], we align visual and language embedding through visual instruction tuning. We leverage a pre-trained video encoder and LLM and then inject the video embedding into LLM through a MLP projector to build a fully differentiable MLLM.
Video Encoder We adopt the pre-trained LanguageBind video encoder [85] as our frozen visual backbone $f_{v}$, which is based on a ViT-B/32 vision transformer [19]. As shown in Fig. 2, given an input video frame sequence $V_{i}=\left\{v_{i}^{1},v_{i}^{2},\ldots,v_{i}^{k}\right\}\in\mathbb{R}^{3\times k\times 224\times 224}$, we first split the video into multiple temporal sequences, each consisting of patches that share the same spatial location across different frames. These patches are then transformed through a linear projection for a vision transformer to output video embedding $\bm{z}_{vo}\in\mathbb{R}^{2048\times 1024}$. The video encoder is pre-trained with video-language contrastive learning (i.e. CLIP4clip [49]) without further fine-tuning.
Figure 2: Video Encoder architecture. Video is first split into $k\times 32\times 32$ patches concatenated in time, where these patches are linear projected to video embedding. Then, the model is trained with video-language contrastive learning (CLIP4clip) [49] to obtain language-align video representation.
@@
In particular, the projector $f_{p}$ takes form in Eq. 1, where we use GELU [26] as an activation function. We train the projector with a two-stage training strategy as detailed in Sec. III-B.
Large Language Model Backbone Finally, the LLM takes in both aligned video embedding $\bm{z}_{v}$ and language embedding $\bm{z}_{t}$ of textual context information and task instruction to predict both textual action explanation and numerical control signal. We adopt Vicuna 1.5 7B [84], which is instruction-tuned based on LLaMA2 [72] as our LLM backbone. For decoder-only LLM conditioned on the multi-modal contextual prefix $\bm{z}_{1:n}=[\bm{z}_{v},\bm{z}_{t}]$ of length $N$, the joint probability of output $\bm{x}_{n+1:L}$ as in Eq. 2, where $P_{\theta}$ is transformer-based LLM backbone parameterized by $\theta$.

### v2

@@ Abstract
We need to trust robots that use often opaque AI methods. They need to explain themselves to us, and we need to trust their explanation. In this regard, explainability plays a critical role in trustworthy autonomous decision-making to foster transparency and acceptance among end users, especially in complex autonomous driving. Recent advancements in Multi-Modal Large Language models (MLLMs) have shown promising potential in enhancing the explainability as a driving agent by producing control predictions along with natural language explanations. However, severe data scarcity due to expensive annotation costs and significant domain gaps between different datasets makes the development of a robust and generalisable system an extremely challenging task. Moreover, the prohibitively expensive training requirements of MLLM and the unsolved problem of catastrophic forgetting further limit their generalisability post-deployment. To address these challenges, we present RAG-Driver, a novel retrieval-augmented multi-modal large language model that leverages in-context learning for high-performance, explainable, and generalisable autonomous driving. By grounding in retrieved expert demonstration, we empirically validate that RAG-Driver achieves state-of-the-art performance in producing driving action explanations, justifications, and control signal prediction. More importantly, it exhibits exceptional zero-shot generalisation capabilities to unseen environments without further training endeavours.
@@ I Introduction
Driven by the emerging development of deep learning, autonomous driving has observed a paradigm shift from rules-based decision systems [schwarting2018planning, faisal2019understanding] to data-driven learning-based approaches [Hu_2023_CVPR, bojarski2016end, kendall2019learning]. However, this comes at the cost of transparency in decision-making, especially for end-to-end autonomous driving systems which are considered black-box in nature [chen2023end]. Thus, in addition to precision in action control, explanation provision is key in ensuring trustworthy decision-making to reconcile the system’s decisions with end-user expectations to foster confidence and acceptance [yang2023survey, bonnefon2016social] in dynamic driving environments.
@@
Traditional approaches have mainly relied on attention visualisation [bojarski2016visualbackprop, bojarski2017explaining, moriVisualExpl] as a proxy to rationalise the decisions of the black-box systems or auxiliary intermediate tasks such as semantic segmentation [hawke2020urban, jain2023ground], object detection [chitta2022transfuser, jaeger2023hidden], and affordance prediction [shao2023safety, li2018rethinking] provide meaningful intermediate representation for decision-making. However, these methods do not engage end-users in the dialogue as they are one-directional and not readily comprehensible by the general users for the purpose of fostering trust and confidence. An alternative promising approach is the integration of natural language explanations [kim2018bddx, jin2023adapt, marcu2023lingoqa], in particular through Multi-Modal Large Language Models (MLLMs) [achiam2023gpt, team2023gemini]. These models, pre-trained on extensive web-scale datasets, demonstrate remarkable reasoning capacity, enabling the transformation of complex vehicular decision-making processes into more understandable narrative formats, thereby offering a new layer of explainability to conventional systems.
@@
While several early attempts have demonstrated the potential of MLLMs as general explainable driving agents [xu2023drivegpt4, wang2023drivemlm, ma2023dolphins], these methods fall short of human-level understanding. One of the limitations is their failure to generalise to unseen environments. A primary obstacle is the lack of high-quality annotated data [nie2023reason2drive], coupled with the significant domain shift across various datasets [domainbed], which hinders the models’ generalisation capacity to novel environments outside of the training data distribution. Another critical challenge is the prohibitively expensive training requirement and the unsolved problem of catastrophic forgetting [kirkpatrick2017overcoming], which make re-training or fine-tuning impractical solutions due to the immense computational demands and severe performance degradation. Consequently, this further limits the models’ generalisability after deployment, as they struggle to effectively utilise new data in constantly evolving environments and driving scenarios.
@@
To address these challenges, we introduce RAG-Driver, a novel retrieval-augmented multi-modal large language model tailored for generalisable and explainable end-to-end driving. As illustrated in Fig.1, it outputs natural language texts corresponding to (1) the driving action and (2) justification of that driving action along with (3) numerical control signals based on driving videos. The natural language texts are aligned with control signals during in-context learning [brown2020language] to enable faithful introspective explanation provision. The novelty of RAG-Driver is the integration of retrieval-augmented in-context learning (RA-ICL) mechanisms that significantly improve generalisation performance in unseen driving environments. It allows efficient recall of similar driving scenarios as contextual information augmenting MLLM prediction through implicit meta-optimisation (Sec. III-C). Through extensive experiments, we show that RAG-Driver outperforms existing methods on both in-domain deployments as well as deployment in unseen environments (without any fine-tuning). By being grounded in analogical demonstrations, our framework significantly reduces the need for continuous retraining while enhancing the generalisability and quality of generated explanatory texts. Our primary contributions are as follows:
@@
    Achieving state-of-the-art introspective driving explanation performance on the standard benchmark BDD-X [kim2018bddx].
@@
    Demonstrating exceptional zero-shot generalisation capacity to unseen scenarios without training effort through a customised dataset Spoken-London featuring video sequences narrated by a professional driving instructor.
@@ II-A Explainable End-to-End Autonomous Driving
End-to-end learned driving [chen2023end] maps directly from raw sensor input to vehicle control signals. This data-driven, joint optimisation of perception, prediction, and planning can be simple and efficient [Hu_2023_CVPR]. In this area, various learning-based approaches, including behavior cloning [bojarski2016end, chen2020learning, pomerleau1988alvinn, zhang2022learning, peng2023learning], inverse optimal control [zeng2019end, sadat2020perceive, wang2021end], and reinforcement learning [kendall2019learning, toromanoff2020end, chekroun2023gri, li2022efficient] are promising. A key area of focus in this area is explainability [zablocki2022explainability], which is crucial for improving transparency and building trust towards wider public acceptance of autonomous systems [marcu2023lingoqa, gadd2020sense, OmeizaExplainSurvey2022]. One line of works leverages attention visualisation – either to directly identify salient regions of input images that are important for driving decision-making [kim2017interpretable, bojarski2016visualbackprop, bojarski2017explaining] or to assist feature aggregation for downstream motion planning tasks [moriVisualExpl, chitta2021neat, xiao2023scaling, renz2022plant]. Another line of work uses intermediate auxiliary tasks such as semantic segmentation [hawke2020urban, jain2023ground], object detection [chitta2022transfuser, jaeger2023hidden], and affordance prediction [shao2023safety, li2018rethinking] which help to decoder latent representations to human-understandable representation. While these methods provide explainable mechanisms by associating decision-making processes with semantic or visual representations, they are not readily comprehensible by the general users for the purpose of fostering trust and confidence.
Alternatively, recent research shows promise in utilising natural language explanation. Several works develop specialist explainers [kim2018bddx, kuhn2023textual] using attention alignment in visual input and textual generation for grounded driving action explanation. ADAPT [jin2023adapt] uses a vision-language transformer with separate decoder for caption generation alongside control signal prediction. More recently, several works explore the potential of Multi-modal Large Language Models (MLLMs). Works such as, DriveGPT4 [xu2023drivegpt4], Lingo [marcu2023lingoqa], and DrivingMLM [wang2023drivemlm] have shown promising potential in general question-answering for driving and action planning. However, a common obstacle in both specialist and MLLM-based generalist models is the scarcity of data due to expensive annotation costs and significant domain gaps between different datasets, making the development of a robust and generalisable model an extremely challenging task. In our work, we aim to overcome these obstacles by employing a more robust inference paradigm of retrieval-augmented in-context learning to bridge domain gaps and circumvent the need for annotations in new domains.
@@ II-B Multi-Modal Large Language Model
Recent advancements in Large Language Models (LLMs) have paved the way for the emergence of Multi-modal Large Language Models (MLLMs) [achiam2023gpt, team2023gemini]. Benefiting from scalable transformer-based architecture and web-scale training data, these models have demonstrated notable capabilities in general-purpose visual understanding tasks. One line of work focuses on modality fusion in the latent space, offering a scalable, end-to-end solution for MLLMs. For instance, Flamingo [alayrac2022flamingo] and BLIP2 [li2023blip] fuse visual tokens to frozen LLM through gated attention and query transformers, respectively. LLaVA [liu2023llava] and MiniGPT4 [zhu2023minigpt] use simple Multi-Layer Perceptron (MLP) with visual instruction tuning to align the pre-trained image encoder to LLM. Most relevant to us is the line of work focusing on video-language models such as Video-LLaVA [lin2023video] and Video-LLaMA [damonlpsg2023videollama], which integrate pre-trained video encoders into LLMs using strategies similar to those in image-based models.
With remarkable perception and reasoning capacity, MLLMs reveal promising potential in various robotics tasks such as reasoning [driess2023palm, ahn2022can, huang2023instruct2act] and planning [rana2023sayplan, brohan2023rt, lin2023text2motion]. The most similar to us is the idea that engages a generalist foundation model for end-to-end embodied agent. PaLM-e [driess2023palm] injects images, state estimates, and other sensor modalities to LLM and autoregressively produces natural language commands. RT-2 [brohan2023rt] and RT-X [padalkar2023open] fine-tune on image and low-level robot control signal pairs to perform end-to-end robot control. Specifically in driving, numerous approaches leverage language-only LLM for decision-making [yang2023survey] and then augment it with external perception module feedback [mao2023language, han2024dme, chen2023driving], designed chain-of-thought reasoning template [mao2023language, ma2023lampilot] or downstream planner [sha2023languagempc, sharan2023llm, chen2023driving] to form a system-level driving agent. Another more relevant line of work is end-to-end driving agents. DriveGPT4 [xu2023drivegpt4] leverages a finetuned video-language model Valley [luo2023valley] on driving specific visual instruction tuning based on BDD-X [kim2018bddx]. Dolphins [ma2023dolphins] further use the designed Grounded chain-of-thought to enhance reasoning capacity. DrivingMLM [wang2023drivemlm] and Reason2Drive [nie2023reason2drive] scale up the driving visual instruction tuning dataset through simulator-based data engine and annotation of existing large-scale datasets, respectively. While these approaches have demonstrated the potentials of MLLM, the prohibitively expensive training cost and the unsolved problem of catastrophic forgetting, which makes re-training or fine-tuning post-deployment challenging, further limit their generalisation capacity to unseen driving environments. To solve this problem, we leverage a training-free retrieval-augmented in-context learning mechanism.
@@ II-C In-Context Learning and Retrieval-Augmented Generation
While LLMs demonstrate strong generative and reasoning capacity, there are still several issues associated with their output, such as hallucination [huang2023survey], and slow knowledge updates [kandpal2023large]. In-context Learning (ICL) [brown2020language, dong2022survey] has emerged as a promising approach in LLM inference, potentially addressing several of these issues. This paradigm involves providing a test query alongside a few demonstration examples as contextual information. The LLM then generates an output for the test instance based on analogies drawn from context, without any updates to its parameters [rubin2021learning]. While ICL has been observed to enhance generalisability in various Natural Language Processing (NLP) tasks, its application in multi-modal contexts remains less explored, potentially due to the challenges associated with curating structured high-quality multi-modal ICL datasets. Retrieval-Augmented Generation (RAG) [lewis2020retrieval] is another important inference paradigm for LLM. It provides an external knowledge database to augment the compressed model knowledge within LLM in inference by dynamically retrieving relevant information pieces as contextual information. One of its promising applications is leveraging a systematic approach to curate In-Context Learning (ICL) examples. In this work, we build upon these inference paradigms and extend their application to Multimodal Large Language Models (MLLMs). We introduce a retrieval-augmented in-context learning mechanism through a curated multi-modal driving in-context instruction tuning dataset and a vector similarity-based retrieval engine specifically tailored for driving applications.
@@ III-A Multi-modal Large Language Model Architecture
Following the successful MLLM paradigm of Video-LLaVA [lin2023video], we align visual and language embedding through visual instruction tuning. We leverage a pre-trained video encoder and LLM and then inject the video embedding into LLM through a MLP projector to build a fully differentiable MLLM.
Video Encoder We adopt the pre-trained LanguageBind video encoder [zhu2023languagebind] as our frozen visual backbone $f_{v}$, which is based on a ViT-B/32 vision transformer [dosovitskiy2020image]. As shown in Fig. 2, given an input video frame sequence $V_{i}=\left\{v_{i}^{1},v_{i}^{2},\ldots,v_{i}^{k}\right\}\in\mathbb{R}^{3\times k\times 224\times 224}$, we first split the video into multiple temporal sequences, each consisting of patches that share the same spatial location across different frames. These patches are then transformed through a linear projection for a vision transformer to output video embedding $\bm{z}_{vo}\in\mathbb{R}^{2048\times 1024}$. The video encoder is pre-trained with video-language contrastive learning (i.e. CLIP4clip [Luo2021CLIP4Clip]) without further fine-tuning.
Figure 2: Video Encoder architecture. Video is first split into $k\times 32\times 32$ patches concatenated in time, where these patches are linear projected to video embedding. Then, the model is trained with video-language contrastive learning (CLIP4clip) [Luo2021CLIP4Clip] to obtain language-align video representation.
@@
In particular, the projector $f_{p}$ takes form in Eq. 1, where we use GELU [hendrycks2016gaussian] as an activation function. We train the projector with a two-stage training strategy as detailed in Sec. III-B.
Large Language Model Backbone Finally, the LLM takes in both aligned video embedding $\bm{z}_{v}$ and language embedding $\bm{z}_{t}$ of textual context information and task instruction to predict both textual action explanation and numerical control signal. We adopt Vicuna 1.5 7B [zheng2023judging], which is instruction-tuned based on LLaMA2 [touvron2023llama2] as our LLM backbone. For decoder-only LLM conditioned on the multi-modal contextual prefix $\bm{z}_{1:n}=[\bm{z}_{v},\bm{z}_{t}]$ of length $N$, the joint probability of output $\bm{x}_{n+1:L}$ as in Eq. 2, where $P_{\theta}$ is transformer-based LLM backbone parameterized by $\theta$.